# Physical properties of circumnuclear ionising clusters. III. Kinematics of gas and stars in NGC 7742

### v2

@@ Abstract
Aims. In this third paper of a series, we study the kinematics of the ionised gas and stars, calculating the dynamical masses of the circumnuclear star-forming regions in the ring of of the face-on spiral NGC 7742.
Methods. We have used high spectral resolution data from the MEGARA instrument attached to the Gran Telescopio Canarias (GTC) to measure the kinematical components of the nebular emission lines of selected HII regions and the stellar velocity dispersions from the CaT absorption lines that allow the derivation of the associated cluster virialized masses.
Results. The emission line profiles show two different kinematical components: a narrow one with velocity dispersion 10 km/s and a broad one with velocity dispersion similar to those found for the stellar absorption lines. The derived star cluster dynamical masses range from 2.5 106 to 10.0 107 M⊙.
@@
Circumnuclear Star Forming Regions (CNSFR), also referred to as “hotspots”, are similar to luminous and large disk HII regions, but look more compact and show higher peak surface brightness (Kennicutt et al., 1989). In early type spirals they are expected to be of higher metallicity as it corresponds to their position near the galactic bulge (Díaz et al., 2007). In many cases, CNSFR contribute substantially to the emission of the entire nuclear region seen in the presence of an active nucleus. In fact, their large H luminosities, typically higher than 1039 erg s-1, overlap with those of HII galaxies pointing to relatively massive star clusters as their ionization source. Hence, these regions constitute excellent places to study how star formation proceeds in circumnuclear environments.
@@
In previous works Hägele et al. (2007, 2009, 2010) analysed long slit high spectral resolution data of CNSFR in three nearby galaxies, NGC 3351, NGC 2903 and NGC 3310. They used the 4.2-m William Herschel Telescope (WHT) and the ISIS spectrograph in the blue and red configurations which provides a comparable velocity resolution of about 13 km/s in both spectral ranges. The H and [OIII]5007 Å emission lines and the CaII triplet lines (CaT) were present in the spectra thus allowing the measurement of the gas and star velocity dispersions which, together with their sizes measured on WPC2-HST camera and assuming virialization, made possible the derivation of the dynamical masses of the young clusters in the regions. These masses, found to be between 4.9 106 and 1.9 108 M⊙, seem to be rather high pointing to star-forming complexes rather than individual clusters. However, the masses of the corresponding ionizing clusters, as derived from H luminosities under the assumption of no photon escape, only accounts for between 1 % and 10 % of the derived dynamical mass which raises the question of the possible presence of more than one stellar population in these massive circumnuclear star clusters.
@@
On the other hand, one of the main results of these works shows that for the three observed galaxies, the gas velocity dispersions as measured from hydrogen recombination lines (H and Paschen emission lines) are found to be considerably smaller (by about 25 km/s) than those measured from the stellar CaT lines. In turn, the velocity dispersion obtained from the [OIII]5007 Å line is very close to the measured stellar one. This result suggests the possible existence of two kinematically distinct components in the gas in all regions, although not all of them being associated with the ionising clusters (see Hägele et al., 2013).
@@
The best possibility to resolve this question is to gather observations with both high spatial and spectral resolution simultaneously. At present, the ideal instrument to do this is MEGARA, a high-resolution integral field spectrograph (IFS) installed on the 10.4-m GTC telescope, covering a field of view (FoV) of 12.5 × 11.3 arcsec2 with a spatial resolution of 0.62 arcsec and a spectral range from 3650 to 9700 Å at three different resolutions, R 6000, 12000, and 20000. The MEGARA spatial and spectral resolution characteristics ensure the proper identification and resolution of the ionising clusters present in CNSFR so that the derived masses obtained can be meaningful and reliable.
@@
In the first paper of this series, we have analysed CNSFRs in the galaxy NGC 7742 using IFS data obtained with the MUSE spectrograph which provides spectral resolutions between R 1770 - 3590 at 4800 and 9300 Å respectively. This galaxy is classified as an SA(r)b galaxy and its morphology is dominated by a circumnuclear ring located at around 1 kpc from the galaxy centre. The galaxy harbours a mildly active nucleus whose spectrum is dominated by emission lines characteristic of both LINER and HII regions. In total, 88 ring HII regions with H fluxes larger than 10-18erg/s/cm2 were selected and characterised using their nebular spectra with the help of photoionization and stellar population synthesis models (Cloudy and PopStar, respectively). These regions seem to be ionised by young clusters with mean ages of about 5 Ma. The chemical composition of the ionised gas was calculated using sulphur as tracer and were found to be between 0.25 and 2.4 times the solar value, with most regions showing values slightly below solar. This result is consistent with the effective temperature of the cluster stars, between 35000 K and 40000 K as estimated from the ratio of HeI to H ionising photons. The high spatial resolution offered by MUSE allowed the measurement of the CNSFR sizes on H flux maps, all of them being spatially resolved. These measured sizes are fully consistent with those predicted by photoionisation models, thus leaving little room for photon escape and justifying the assumption of the regions being radiation bounded. The young ionising star clusters powering this radiation have masses around 3.5 104 M⊙, comparable to the mass of ionised gas and about 19 % of the corrected photometric cluster mass that includes a young non ionising stellar population whose estimated age is around 300 Ma. A minor merger scenario is proposed for the formation of these clusters due to the counter rotating nature of the ring (Pizzella et al., 2004; de Zeeuw et al., 2002) and the homogeneity of abundances and continuum colours.
@@
The present work is focused on three principal objectives. First, testing the capabilities of the new MEGARA instrument for the study of resolved kinematical components in the ionised gas of HII regions; second, determining the kinematic origin of such components if present, for which observing both recombination lines (H) and collisionally excited lines ([NII] and [SIII] is needed; and third, measuring the stellar velocity dispersion along the line of sight from the CaT lines in order to derive cluster dynamical masses. Accomplishing these objectives will ensure the cluster origin of the CaT lines and the reliability of this derivation. It should be said that a proper comparison between ionising and dynamical masses for ionising young star clusters has not yet been done and is much needed.
@@
In Section 2, we describe the observations and the reduction of the data, focusing on the new MEGARA observations. Section 3 is devoted to the presentation of our results. The discussion of cluster sizes and dynamical masses and the comparison between gas and star kinematical components is presented in Section 4. Finally, Section 5 summarizes this work and our conclusions.
@@ 2 Observations and data reduction
In this work, we have used data for the ring HII regions of NGC 7742 (J2000, RA=356.065542 deg, DEC=10.767083 deg, Skrutskie et al., 2006) obtained with the high resolution integral field spectrograph MEGARA (Multi-Espectrógrafo en GTC de Alta Resolución para Astronomía, Carrasco et al., 2018) attached to the GTC 10.4 m telescope at La Palma Observatory. We have used MEGARA in the integral-field Unit (IFU) mode working in the visible wavelength range from 3655 to 9700 Å and providing a high spectral (up to RFWHM = 20000) and spatial resolution (0.62 arcsec spaxel size) simultaneously in a moderate size field of view (FoV) (12.5 11.3 arcsec2). It has a total of 623 hexagonal fibers distributed in 567 science-fibers and 56 sky-fibers located at the edges of the FoV at distances from 1.7 arcmin to 2.5 arcmin from the centre of the fiber bundle.
The observations were acquired during 3 nights on 2021 August 28th, 31th and September 14th111Proposal GTC28-21A: A high spatial and spectral resolution study of circumnuclear star-forming regions
PI: Ángeles I. Díaz. Semester 2021-A. under spectroscopic conditions. The seeing during the observations ranged between approximately 0.7 and 0.9 arcsec and the airmass of the science images was always less than 1.25. Four pointings were needed to cover the whole circumnuclear ring extension of NGC 7742 (approximately the central 24x24 arcsec2, see Fig. 1).
@@
a From North toward East. The orientation of 0 deg implies that the North is aligned with the short side and the East-West axis with the long side of the detector.
b Total target observed time without overheads.
Three dispersion elements (VPHs) were used: VPH863_HR (HR-I, RFWHM 20500), including the calcium triplet stellar lines (CaT) at high spectral resolution; VPH890_LR (LR-Z, RFWHM 5800) including the [SIII]9069,9532 Å emission lines and the CaT lines at low resolution; and VPH665_HR (HR-R, RFWHM 20050) including the [NII]6548,84 Å, H, and [SII]6717,32 Å emission lines at high resolution. A relation of the total integration times and their splitting in different exposures for each of the used setups is given in Tab. 1. Using these configurations, velocity dispersions around 6.4 km/s and 22.5 km/s in FWHM were expected at high and low spectral resolutions respectively.
Calibration images were acquired on the observing nights and on the preceding and following ones. They consisted of: bias frames to create a CCD spatial response function; flat field frames necessary to locate each fiber spectrum on the CCD and to perform a spatial instrumental response function; arc calibration exposures needed for wavelength calibration; and 3 spectrophotometric standard star exposures needed for performing a spectral instrumental response function. The standard procedures of: bias subtraction; tracing and extraction of each fiber spectra; flat field correction; and wavelength calibration were followed applying some required modifications to the standard MEGARA pipeline. (Pascual et al., 2018) 222see https://github.com/guaix-ucm/megaradrp/ using calibration files present in Cardiel & Pascual (2018). Flux calibration was not needed for our work and we decided not to do it in order not to introduce additional errors in our subsequent analysis. After the reduction procedure, a data cube per pointing was produced (i. e. 8, 5, and 4 cubes for the HR-R, HR-I and LR-Z setups respectively) with a spatial binning of 0.2 arcsec/pix.
@@
All wavelengths are in the rest frame.
Astrometric aligning was achieved by comparing our MEGARA data with those extracted from the publicly available observations of NGC 7742 obtained by the IFS MUSE (Science Verification run on 2014 June 22, ESO Programme 60.A-9301(A), M. Sarzi) and already analysed in Zamora & Díaz (2023b). The 2D cross correlation technique proposed by Tonry & Davis (1979) was used to find the spatial shift between MEGARA and MUSE data. We have constructed H and [SIII]9532 Å emission line maps for the HR-R and LR-Z setups respectively, and a continuum map near the CaT absorption lines for the HR-I setup. For the emission line maps, a linear behavior of the continuum emission in the region of interest was assumed choosing appropriate side-bands of a given width for each line. Table 2 lists in columns 1 to 5: (1) the identification of each line; (2) its central wavelength; (3) its width; and (4,5) the limits of the two continuum side-bands. The upper and lower table panels show these parameters for MEGARA and MUSE data respectively. The followed procedure ensures that our pointings are correctly aligned.
Finally, we have combined the overlapping individual cubes in a single one using the average of the data available in each pixel with no spatial smoothing applied. Figure 2 shows the resulting cube for HR-R setup as an example. It is compressed in the spectral direction and superimposed are contours of data obtained with MUSE showing the consistency with those obtained from MEGARA.
@@ 3.1 HII region selection
The HII region selection has been accomplished as described in the first paper of this series (see Zamora & Díaz, 2023b) consisting in the application of an iterative procedure for the detection of high intensity clumps on an H emission line map. This method requires as input the maximum and minimum extent of regions (500 pc and the point spread function of observations respectively) and the relative flux intensity in each of the regions with respect to the emission of its centre (10%). Then, the program tried different values for the absolute flux intensity background of the complete map adopting the one that minimizes the dispersion of the spatial residuals in the map.
We consider that this HII region segmentation is appropriate for our MEGARA study since the astrometry correction has been made with respect to MUSE data. Also, since the selected pixel scale for the cube construction is the same as the MUSE one (0.2 arcsec/pix) we can assure that all the ring HII regions are spatially resolved. Additionally, the measured angular radii of the regions calculated using this method were found to be compatible with radiation bounded ionised nebulae (see Zamora & Díaz, 2023b). Regions R2, R13, R24, R39, R40, R46, R49, R68, and R82 fall outside the area covered by the FoV of MEGARA, and therefore are not included in the present study. In the end, we have selected a total of 79 HII regions in the ring.
@@ 3.2 Nebular emission line measurements
We have extracted each region spectrum by integrating the flux inside its corresponding aperture and we have used the code LiMe (LIne MEasuring library Fernández et al., 2022) to separate the different kinematical components present in each emission line. This program uses a defined window for each line and two additional ones to calculate the continuum below. After the fitting, we have taken into account only those components that meet the requirement: , with being the Gaussian amplitude and the local standard deviation of the continuum bands selected for each fit. Next, we have measured the radial velocity, , and the velocity dispersion, , of each component in the following emission lines: H, [NII]6548,84 Å and [SII]6716,31 Å from the HR-R setup, and [SIII]9069,9532 Å from the LR-Z ones. We have performed an individual fitting for each emission line and adopted a kinematical bond between the components of lines of the same element: [NII]6548 Å with [NII]6584 Å; [SII]6716 Å with [SII]6731 Å; and [SIII]9069 Å with [SIII]9532 Å.
Fig. 3 shows two examples of the procedure for regions R1 and R9, showing in different panels from left to right the emission lines: H and [NII]6548,84 Å; [SII]6716,31 Å; and [SIII]9069 Å and [SIII]9532 Å. Eight regions out of 79 ( 10%), labeled R6, R48, R78, R79, R81, R84, R85, and R88, have been found to have one only component; the rest of them show at least two in the H emission line; and only one region, R1, shows two components in the [SIII] lines. In fact, in this region, which is shown in the upper panel of the figure, three components are clearly identified in the H and [NII] lines, and also a third can be appreciable in the[SIII]9532 Å emission line. The region is located near a supernova event: SN 1993R.
@@
Measured velocities have been corrected to rest frame using the redshift derived value of v = 1652.364 km/s (z = 0.00553). Errors in the velocity of each component are given by the LiMe code using a Monte-Carlo procedure and the noise calculated in the continuum bands. Table 3 shows the velocity results listing in columns 1 to 6: (1) the region ID; (2) the line component; and (3-6) the velocity of the H, [NII], [SII], and [SIII] emission lines respectively. The corresponding errors in the velocity dispersions have been calculated as the difference between the FWHM of the largest and smallest fitted Gaussian function taking into account the noise of the fitted spectrum. Using the amplitude of the Gaussian fitted, , and the noise calculated in the residuals of the fit, , the error can be calculated as:
The measured velocity dispersions of the gas needs to be corrected for instrumental dispersion, and thermal broadening, , hence the corrected value for the gas velocity dispersion will be:
The instrumental dispersion values are 6.4 km/s and 22.5 km/s for the MEGARA high and low resolution configurations respectively (see Sec. 2). The thermal broadening can be calculated assuming a Maxwellian velocity distribution of the ions using the following equation:
where k is Boltzmann’ constant (k=1.380649 10-23 J/K), te is the electron temperature in degrees Kelvin and m is the mass of the ion involed (m = 1.6726 10-27 kg; m = 2.3257 10-26 kg; m = 5.3244 10-26 kg; m = 5.3243 10-26 kg). We have assumed the electron temperature of [SIII] (te[SIII]) calculated in Zamora & Díaz (2023b) for each CNSFR, and its average value is 7967 K, hence the mean thermal corrections are 8.11 km/s for H+; 2.17 km/s for N+; and 1.44 km/s for S+ and S++.

### v3

# Physical properties of circumnuclear ionising clusters III. Kinematics of gas and stars in NGC 7742

@@ Abstract
Aims. For this third paper in the series, we studied the kinematics of the ionised gas and stars, and calculated the dynamical masses of the circumnuclear star-forming regions in the ring of the face-on spiral NGC 7742.
Methods. We used high spectral resolution data from the MEGARA instrument mounted on the Gran Telescopio Canarias (GTC) to measure the kinematical components of the nebular emission lines of selected HII regions and the stellar velocity dispersions from the CaT absorption lines that allow the derivation of the associated cluster virialised masses.
Results. The emission line profiles show two different kinematical components: a narrow one with a velocity dispersion of 10 km/s and a broad one with a velocity dispersion similar to the values found for the stellar absorption lines. The derived star cluster dynamical masses range from 2.5 106 to 10.0 107 M⊙.
@@
Observations of central regions have revealed that nuclear rings are present in at least 20 % of spiral galaxies, with this percentage increasing in galaxies hosting an active galactic nucleus (AGN) (see Knapen, 2005). These structures, typically with diameters of hundreds of parsecs, exhibit high star formation rates that have been studied by several authors over the past decades across different wavelengths (e.g. Levy et al., 2022; Lopez et al., 2023; Brandl et al., 2012; Böker et al., 2008; Mazzuca et al., 2008). These star-forming regions, often referred to as hotspots or circumnuclear star-forming regions (CNSFRs), are similar to luminous and large disk HII regions, but look more compact and show higher peak surface brightness (Kennicutt et al., 1989). In early-type spirals they are expected to be of higher metallicity as it corresponds to their position near the galactic bulge (Díaz et al., 2007). In many cases, CNSFRs contribute substantially to the emission of the entire nuclear region seen in the presence of an active nucleus. Their high H luminosities, typically higher than 1039 erg s-1, overlap with those of HII galaxies, and point to relatively massive star clusters as their ionisation source. Hence, these regions are excellent places to study how star formation proceeds in circumnuclear environments; they are crucial for understanding the connection between nuclear activity and the star formation processes occurring in these hostile environments (see Sarzi et al., 2007).
@@
In previous works, Hägele et al. (2007, 2009, 2010) analysed long-slit high spectral resolution data of CNSFRs in three nearby galaxies: NGC 3351, NGC 2903, and NGC 3310. They used the 4.2m William Herschel Telescope (WHT) and the ISIS spectrograph in the blue and red configurations, which provides a comparable velocity resolution of about 13 km/s in both spectral ranges. The H and [OIII]5007 Å emission lines and the CaII triplet lines (CaT) were present in the spectra, thus allowing the measurement of the gas and star velocity dispersions, which, together with their sizes measured on the WPC2-HST camera and assuming virialisation, made possible the derivation of the dynamical masses of the young clusters in the regions. These masses, found to be between 4.9 106 and 1.9 108 M⊙, seem to be rather high and suggest star-forming complexes rather than individual clusters. However, the masses of the corresponding ionising clusters, as derived from H luminosities under the assumption of no photon escape, only accounts for between 1 % and 10 % of the derived dynamical mass, which raises the question of the possible presence of more than one stellar population in these massive circumnuclear star clusters.
@@
On the other hand, one of the main results of these works shows that for the three observed galaxies, the gas velocity dispersions as measured from hydrogen recombination lines (H and Paschen emission lines) are found to be considerably smaller (by about 25 km/s) than those measured from the stellar CaT lines. In turn, the velocity dispersion obtained from the [OIII]5007 Å line is very close to the measured stellar value. This result suggests the possible existence of two kinematically distinct components in the gas in all regions, although not all of them are associated with the ionising clusters (see Hägele et al., 2013).
@@
The best possibility to resolve this question is to gather observations with both high spatial and spectral resolution simultaneously. At present, the ideal instrument for this is MEGARA, a high-resolution integral field spectrograph (IFS) installed on the 10.4m GTC telescope, covering a field of view (FoV) of 12.5 × 11.3 arcsec2 with a spatial resolution of 0.62 arcsec and a spectral range from 3650 to 9700 Å at three different resolutions, R 6000, 12000, and 20000. The MEGARA spatial and spectral resolution characteristics ensure the proper identification and resolution of the ionising clusters present in CNSFRs so that the derived masses obtained can be meaningful and reliable.
@@
In the first paper of this series we analyse CNSFRs in the galaxy NGC 7742 using IFS data obtained with the MUSE spectrograph, which provides spectral resolutions between R 1770 - 3590 at 4800 and 9300 Å, respectively. This galaxy is classified as an SA(r)b galaxy and its morphology is dominated by a circumnuclear ring located at around 1 kpc from the galaxy centre. The galaxy harbours a mildly active nucleus whose spectrum is dominated by emission lines characteristic of both LINER and HII regions. In total, 88 ring HII regions with H fluxes larger than 10-18erg/s/cm2 were selected and characterised using their nebular spectra with the help of photoionisation and stellar population synthesis models (Cloudy and PopStar, respectively). These regions seem to be ionised by young clusters with mean ages of about 5 Ma. The chemical composition of the ionised gas was calculated using sulphur as a tracer and were found to be between 0.25 and 2.4 times the solar value, with most regions showing values slightly below solar. This result is consistent with an effective temperature of the cluster stars between 35000 K and 40000 K, as estimated from the ratio of HeI to H ionising photons. The high spatial resolution offered by MUSE allowed the measurement of the CNSFR sizes on the H flux maps, all of which were spatially resolved. These measured sizes are fully consistent with those predicted by photoionisation models, thus leaving little room for photon escape and justifying the assumption of the regions being radiation bounded. The young ionising star clusters powering this radiation have masses of around 3.5 104 M⊙, comparable to the mass of ionised gas and about 19 % of the corrected photometric cluster mass that includes a young non-ionising stellar population whose estimated age is around 300 Ma. A minor merger scenario is proposed for the formation of these clusters due to the counter rotating nature of the ring (Pizzella et al., 2004; de Zeeuw et al., 2002) and the homogeneity of abundances and continuum colours.
@@
The present work is focused on three principal objectives: first, testing the capabilities of the new MEGARA instrument for the study of resolved kinematical components in the ionised gas of HII regions; second, determining the kinematic origin of such components if present, for which observing both recombination lines (H) and collisionally excited lines ([NII] and [SIII] is needed; and third, measuring the stellar velocity dispersion along the line of sight from the CaT lines in order to derive cluster dynamical masses. Accomplishing these objectives will ensure the cluster origin of the CaT lines and the reliability of this derivation. It should be said that a proper comparison between ionising and dynamical masses for ionising young star clusters has not yet been done and is much needed.
@@
In Sect. 2 we describe the observations and the reduction of the data, focusing on the new MEGARA observations. Section 3 is devoted to the presentation of our results. The discussion of cluster sizes and dynamical masses and the comparison between gas and star kinematical components is presented in Sect. 4. Finally, in Sect. 5 we summarise our work and present our conclusions.
@@ 2 Observations and data reduction
For this work we used data for the ring HII regions of NGC 7742 (J2000, RA=356.065542 deg, DEC=10.767083 deg, Skrutskie et al., 2006) obtained with the high-resolution integral field spectrograph MEGARA (Multi-Espectrógrafo en GTC de Alta Resolución para Astronomía, Carrasco et al., 2018) attached to the GTC 10.4 m telescope at La Palma Observatory. We used MEGARA in the integral-field Unit (IFU) mode working in the visible wavelength range from 3655 to 9700 Å and providing a high spectral resolution (up to RFWHM = 20000) and spatial resolution (0.62 arcsec spaxel size) simultaneously in a moderate size field of view (FoV) (12.5 11.3 arcsec2). It has a total of 623 hexagonal fibres distributed in 567 science-fibres and 56 sky-fibres located at the edges of the FoV at distances from 1.7 arcmin to 2.5 arcmin from the centre of the fibre bundle.
The observations were acquired during three nights on 2021 August 28 and 31, and September 14111Proposal GTC28-21A: A high spatial and spectral resolution study of circumnuclear star-forming regions
PI: Ángeles I. Díaz. Semester 2021-A. under spectroscopic conditions. The seeing during the observations ranged between approximately 0.7 and 0.9 arcsec and the airmass of the science images was always less than 1.25. Four pointings were needed to cover the whole circumnuclear ring extension of NGC 7742 (approximately the central 24x24 arcsec2; see Fig. 1).
@@
Three dispersion elements (VPHs) were used: VPH863_HR (HR-I, RFWHM 20500), including the calcium triplet stellar lines (CaT) at high spectral resolution; VPH890_LR (LR-Z, RFWHM 5800), including the [SIII]9069,9532 Å emission lines and the CaT lines at low resolution; and VPH665_HR (HR-R, RFWHM 20050), including the [NII]6548,84 Å, H, and [SII]6717,32 Å emission lines at high resolution. A relation of the total integration times and their splitting in different exposures for each of the used set-ups is given in Table 1. Using these configurations, velocity dispersions of around 6.4 km/s and 22.5 km/s in full width at half maximum (FWHM) were expected at high and low spectral resolutions, respectively.
Calibration images were acquired on the observing nights and on the preceding and following nights. They consisted of bias frames to create a CCD spatial response function; flat-field frames necessary for locating each fibre spectrum on the CCD and for performing a spatial instrumental response function; arc calibration exposures needed for wavelength calibration; and three spectrophotometric standard star exposures needed for performing a spectral instrumental response function. The standard procedures of bias subtraction, tracing and extraction of each fibre spectra, flat-field correction, and wavelength calibration were followed applying some required modifications to the standard MEGARA pipeline (Pascual et al., 2018)333see https://github.com/guaix-ucm/megaradrp/ using the calibration files present in Cardiel & Pascual (2018). Flux calibration was not needed for our work, and we decided not to do it in order not to introduce additional errors in our subsequent analysis. After the reduction procedure, one data cube per pointing was produced (i.e. eight, five, and four cubes for the HR-R, HR-I, and LR-Z set-ups, respectively) with a spatial binning of 0.2 arcsec/pix.
@@
Astrometric alignment was achieved by comparing our MEGARA data with those extracted from the publicly available observations of NGC 7742 obtained by the IFS MUSE (Science Verification run on 2014 June 22, ESO Programme 60.A-9301(A), M. Sarzi) and already analysed in Zamora & Díaz (2023b). The 2D cross-correlation technique proposed by Tonry & Davis (1979) was used to find the spatial shift between MEGARA and MUSE data. We constructed H and [SIII]9532 Å emission line maps for the HR-R and LR-Z set-ups, respectively, and a continuum map near the CaT absorption lines for the HR-I set-up. For the emission line maps, a linear behaviour of the continuum emission in the region of interest was assumed choosing appropriate side-bands of a given width for each line. Table 2 lists in Cols. 1 to 5 (from left to right): the identification of each line; its central wavelength; its width; and the limits of the two continuum side-bands. The upper and lower table panels show these parameters for the MEGARA and MUSE data, respectively. The followed procedure ensures that our pointings are correctly aligned.
Finally, we combined the overlapping individual cubes into a single cube using the average of the data available in each pixel with no spatial smoothing applied. Figure 2 shows the resulting cube for HR-R set-up as an example. It is compressed in the spectral direction, and the contours of the data obtained with MUSE showing the consistency with those obtained from MEGARA are superimposed.
@@ 3.1 HII region selection
The HII region selection was accomplished as described in the first paper of this series (see Zamora & Díaz, 2023b) consisting in the application of an iterative procedure for the detection of high-intensity clumps on an H emission line map. This method requires as input the maximum and minimum extent of regions (500 pc and the point spread function of observations, respectively) and the relative flux intensity in each of the regions with respect to the emission of its centre (10%). Then, the program tried different values for the absolute flux intensity background of the complete map adopting the one that minimises the dispersion of the spatial residuals in the map.
We consider that this HII region segmentation is appropriate for our MEGARA study since the astrometry correction was made with respect to MUSE data. In addition, since the selected pixel scale for the cube construction is the same as the MUSE scale (0.2 arcsec/pix) we can assure that all the ring HII regions are spatially resolved. Additionally, the measured angular radii of the regions calculated using this method were found to be compatible with radiation bounded ionised nebulae (see Zamora & Díaz, 2023b). Regions R2, R13, R24, R39, R40, R46, R49, R68, and R82 fall outside the area covered by the FoV of MEGARA, and therefore are not included in the present study. In the end, we selected a total of 79 HII regions in the ring.
@@ 3.2 Nebular emission line measurements
We extracted each region spectrum by integrating the flux inside its corresponding aperture and we used the code LiMe (LIne MEasuring library Fernández et al., 2022) to separate the different kinematical components present in each emission line. This program uses a defined window for each line and two additional ones to calculate the continuum below. After the fitting, we took into account only those components that meet the requirement , with being the Gaussian amplitude and the local standard deviation of the continuum bands selected for each fit. Next, we measured the radial velocity, , and the velocity dispersion, , of each component in the following emission lines: H, [NII]6548,84 Å and [SII]6716,31 Å from the HR-R set-up, and [SIII]9069,9532 Å from the LR-Z set-up. We performed an individual fitting for each emission line and adopted a kinematical bond between the components of lines of the same element: [NII]6548 Å with [NII]6584 Å; [SII]6716 Å with [SII]6731 Å; and [SIII]9069 Å with [SIII]9532 Å.
Figure 3 shows two examples of the procedure for regions R1 and R9; the panels show (from left to right) the following emission lines: H and [NII]6548,84 Å; [SII]6716,31 Å; and [SIII]9069 Å and [SIII]9532 Å. Eight regions out of 79 ( 10%), labelled R6, R48, R78, R79, R81, R84, R85, and R88, were found to have one only component; the rest show at least two components in the H emission line; and only one region, R1, shows two components in the [SIII] lines. In this region, which is shown in the top panel of the figure, three components are clearly identified in the H and [NII] lines, and also a second can be seen in the[SIII]9532 Å emission line. The region is located near a supernova event: SN 1993R.
@@
The measured velocities were corrected to rest frame using the redshift derived value of v = 1652.364 km/s (z = 0.00553). Errors in the velocity of each component are given by the LiMe code using a Monte Carlo procedure and the noise calculated in the continuum bands. Table 3666The accuracy of the fit and the number of components have been determined by analysing the chi-square of the residuals. shows the velocity results in columns 1 to 6: (1) the region ID; (2) the line component; and (3-6) the velocity of the H, [NII], [SII], and [SIII] emission lines, respectively. The corresponding errors in the velocity dispersions were calculated as the difference between the FWHM of the largest and smallest fitted Gaussian function taking into account the noise of the fitted spectrum. Using the amplitude of the fitted Gaussian, , and the noise calculated in the residuals of the fit, , the error can be calculated as
The measured velocity dispersions of the gas needs to be corrected for instrumental dispersion, and thermal broadening, , and hence the corrected value for the gas velocity dispersion is
The instrumental dispersion values are 6.4 km/s and 22.5 km/s for the MEGARA high- and low-resolution configurations, respectively (see Sect. 2). The thermal broadening can be calculated assuming a Maxwellian velocity distribution of the ions using the equation
where k is the Boltzmann constant (k=1.380649 10-23 J/K), te is the electron temperature in degrees Kelvin, and m is the mass of the ion involved (m = 1.6726 10-27 kg; m = 2.3257 10-26 kg; m = 5.3244 10-26 kg; m = 5.3243 10-26 kg). We assumed the electron temperature of [SIII] (te[SIII]) calculated in Zamora & Díaz (2023b) for each CNSFR; its average value is 7967 K, and hence the mean thermal corrections are 8.11 km/s for H+; 2.17 km/s for N+; and 1.44 km/s for S+ and S++.